OpenStudy (anonymous):

What is the axis of symmetry for the function? y = x2 + 3x - 4 A. y = -2 B. y = -1.5 C. x = -2 D. x = -1.5

OpenStudy (anonymous):

is it b

OpenStudy (anonymous):

the axis of symmetry can be found by plugging in the correct values in: -b/2ac

OpenStudy (turingtest):

D. x = -1.5

OpenStudy (turingtest):

I don't think they care how to do the problem.

OpenStudy (anonymous):

i mean -b/2a

OpenStudy (anonymous):

vertex is at -b/2a = -3/2 = -1.5 x = -1.5 D